# Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning

### v1

@@ 4.2 Finetuning Experimental Settings
Datasets. During the fine-tuning, to demonstrate the effectiveness for both data-efficient learning and open-world recognition of the proposed WS3D and WS3D++ under the limited scene reconstruction labeling scheme, we have tested it on various benchmarks, including S3DIS [21], ScanNet [129], and SemanticKITTI [18] for semantic segmentation, and ScanNet [129] for instance segmentation, respectively. Detailed information on each dataset and training details are put into the Appendix.
@@
TABLE I: Comparison of semantic segmentation results with different labeling percentages on ScanNet validation set, S3DIS validation set (Area 5), and SemanticKITTI validation set (Sequence 08). ’Sup-only-GPC’ denotes GPC model trained with only labeled data. ’WS3D’ denotes model trained with our proposed methods. We have shown the performance increase in the last row for each dataset, compared to merely trained models with labeled data (left value) and to the SOTAs GPC [121] (right value).
@@
ScanNet Sup-only-GPC 40.9 48.1 57.2 61.3 64.0 65.3 67.1 68.8 72.9
@@
WS3D [4] 49.9 +9.0 56.2 +8.1 62.2 +5.0 65.8 +4.5 68.5 +4.5 69.4 +4.1 70.3 +3.2 73.4 +4.6 76.9 +4.0
WS3D++ 52.6 +11.7 59.1 +11.0 65.2 +8.0 66.9 +5.6 71.8 +6.6 71.1 +5.8 72.6 +5.5 75.3 +6.6 81.7 +8.8
S3DIS Sup-only-GPC 36.3 45.0 52.9 53.8 59.9 60.3 61.2 62.6 66.4
@@
WS3D [4] 45.3 +9.0 54.6 +9.6 59.3 +6.4 62.3 +7.0 65.7 +5.8 66.5 +6.8 67.2 +6.0 69.5 +6.9 72.9 +6.5
WS3D++ 48.6 +12.3 57.7 +12.7 61.2 +8.3 66.9 +11.6 70.6 +10.7 71.1 +10.8 72.6 +11.4 75.3 +12.7 81.7 +15.3
SemanticKITTI Sup-only-GPC 28.6 34.8 43.9 47.9 53.8 55.1 55.4 57.4 65.0
@@
WS3D [4] 38.9 +10.3 43.7 +8.9 52.3 +8.4 55.5 +7.6 61.4 +7.6 61.8 +6.7 62.1 +6.7 63.2 +5.8 66.9 +1.9
WS3D++ 46.8 +18.2 58.6 +13.8 55.2 +11.3 62.8 +14.9 65.9 +12.1 67.9 +12.8 68.6 +13.2 70.9 +13.5 76.8 +11.8
Nuscene Sup-only-GPC 39.8 45.9 47.1 48.6 55.7 56.3 58.9 59.8 67.9
PRKT [135] 48.0 49.3 51.8 53.5 56.7 59.6 63.8 66.7 70.1
SLidR [134] 48.2 50.3 52.9 55.8 56.9 59.1 60.8 63.8 66.9
PointContrast [27] 48.3 50.2 55.1 58.1 60.2 61.1 62.1 64.9 69.2
CLIP2Scene [27] 56.3 56.8 58.7 61.2 63.6 64.1 64.5 64.8 65.1
LESS [134] 49.1 42.5 50.5 53.9 59.5 59.6 60.5 63.5 66.9
Mix3D [130] 37.7 42.9 50.8 54.1 59.3 60.2 61.8 61.3 68.8
WS3D [4] 49.1 +9.3 51.6 +5.7 52.3 +5.2 55.5 +6.9 61.4 +6.7 61.8 +5.5 62.1 +3.2 63.9 +4.1 71.6 +3.7
WS3D++ [4] 53.7 +13.9 56.7 +10.8 55.2 +11.3 62.8 +14.9 71.6 +15.9 67.9 +11.6 68.6 +9.7 71.9 +12.1 77.8 +9.9
@@
[…]
@@ 4.3 Data-efficient 3D Semantic Segmentation
Overall Experimental Results. For the semantic segmentation, we have tested WS3D on versatile indoor and outdoor benchmarks, including ScanNet [20], S3DIS [21], and SemanticKITTI [18]. We have done extensive experiments with limited labeled data, e.g., only {1%, 5%, 10%, 15%, 20%, 25%, 30%, 40%, 100%} data in training set are available as labeled data. The qualitative results are shown in Fig. 4. In the meanwhile, the quantitative semantic segmentation performance is summarized in Table I. As mentioned, we have used the voxel-based method SparseConv [59] as the backbone. Our WSL model significantly surpasses the supervised-only model in GPC that is merely trained with labeled data, showing that our WSL can effectively make use of the unlabeled data to enhance the feature discrimination capacity of the model. Also, it can be observed the increment of performance is more obvious when the unlabeled data percentage is larger. For example, the performance increase on SemanticKITTI is 10.3% for the 1% labeling percentage, 5.8% for the 40% labeling percentage, and 1.9% for the 100% labeling percentage. This can be possibly explained by the fact that for more unlabeled data, our proposed WS3D can extract more meaningful semantic information from the unlabeled data based on our boundary-guided energy-based loss and confidence-guided region-level contrastive learning design. In addition, compared with current SOTA GPC, our proposed WS3D also achieves consistently better results in semantic segmentation performance, especially when faced with very limited label circumstances (e.g. 1% labeling points). In that case, WS3D outperforms GPC by 3.3%, 7.1%, and 4.2% for ScanNet, S3DIS, and SemanticKITTI, respectively. Fig. 4 shows that we can provide comparable performance compared with fully supervised SOTAs BAAF-Net [89] and Cylinder3D [119] on SemanticKITTI with 5% labels. As shown in Table I, the performance of our enhanced approach WS3D++ has remarkably increased performance compared with WS3D and previous SOTAs, demonstrating the effectiveness of our proposed vision-language knowledge-associated pre-training.
Figure 11: Segmentation result visualizations and comparisons with CLIP text encoder prompts for the outdoor KITTI benchmark. It can be demonstrated the foreground object awareness can be clearly captured comared with the previous SOTAs PLA [85]. Meanwhile, as shown in the last three columns, we can provide clear segmentation for the corresponding visual objects based on the textual prompts.
@@
TABLE II: Comparison of experimental results on 20% and fully labeled case for the task of inductive and transductive learning for our proposed WS3D++, respectively. In transductive learning, the test set is also utilized for network training. We test the task of semantic segmentation on ScanNet, S3DIS, and SemanticKITTI with the evaluation metric of mIoU(%).
@@
ScanNetv2-WS3D 64.0 68.5 71.4 72.9 76.9 77.6
ScanNetv2-WS3D++ 67.3 73.9 75.6 77.3 80.3 81.6
S3DIS Area5 Val. 59.9 65.7 66.6 66.4 72.9 73.5
S3DIS WS3D 61.3 67.6 68.6 68.2 74.8 75.8
S3DIS WS3D++ 62.8 68.9 69.9 69.7 75.5 76.9
Semantic KITTI Val. 53.8 61.4 64.5 65.0 66.9 68.2
Semantic KITTI Val. WS3D++ 54.7 62.9 66.2 67.1 68.2 69.7
Semantic KITTI Test. 55.7 62.5 63.6 65.4 68.1 71.3
Semantic KITTI Test. WS3D++ 56.6 63.7 65.3 67.1 69.7 72.8
@@
TABLE III: Comparison of current State-of-the-art (SOTA) approaches in the limited reconstruction case for the 3D object detection tasks with different ratios of labeled data. The mean average precision (mAP) given by the mean $\pm$ standard deviation in three runs of diverse random splits are reported.
Datasets Models 5% 10% 20%
Induct. Transduct. Induct. Transduct. Induct. Transduct.
SUN RGB-D [137, 66] Baseline 29.9 $\pm$ 1.5 33.5 $\pm$ 0.8 34.4$\pm$ 1.1 40.7$\pm$ 0.9 41.1 $\pm$ 0.3 47.5 $\pm$ 0.5
SESS [138] 34.2 $\pm$ 2.0 38.1 $\pm$ 0.7 42.9 $\pm$ 0.8 45.3$\pm$ 0.9 47.9 $\pm$ 0.4 51.6 $\pm$ 0.3
3D-IOUMatch [139] 39.1 $\pm$ 1.9 46.3 $\pm$ 0.7 45.5 $\pm$ 1.5 53.5 $\pm$ 0.3 49.7 $\pm$ 0.5 54.3 $\pm$ 0.8
SPD [140] 38.5 $\pm$ 0.7 44.3 $\pm$ 0.8 46.0 $\pm$ 1.0 51.7 $\pm$ 0.9 49.6 $\pm$ 0.5 54.5 $\pm$ 0.9
WS3D (Ours) 46.9 $\pm$ 0.8 49.5 $\pm$ 0.7 47.8 $\pm$ 0.8 55.7 $\pm$ 0.8 55.6 $\pm$ 0.7 57.9 $\pm$ 1.2
WS3D-Open (Ours) 47.8 $\pm$ 0.8 50.3 $\pm$ 0.6 49.3 $\pm$ 0.7 58.2 $\pm$ 0.7 59.1 $\pm$ 0.8 60.3 $\pm$ 1.1
WS3D++ (Ours) 52.7 $\pm$ 0.7 54.5 $\pm$ 0.8 53.7 $\pm$ 0.9 63.8 $\pm$ 0.8 65.8 $\pm$ 0.7 65.6 $\pm$ 1.1
ScanNet-V2 [20] Baseline 27.9 $\pm$ 0.5 30.8 $\pm$ 1.5 31.1 $\pm$ 0.7 33.2 $\pm$ 0.5 41.6 $\pm$ 0.9 44.5 $\pm$ 1.2
SESS [138] 32.2 $\pm$ 0.8 36.8 $\pm$ 1.1 39.7 $\pm$ 0.9 44.5 $\pm$ 0.5 47.9 $\pm$ 0.4 49.2 $\pm$ 0.7
3D-IOUMatch [139] 40.0 $\pm$ 0.9 46.3 $\pm$ 0.7 47.2 $\pm$ 0.4 49.6 $\pm$ 0.6 52.8 $\pm$ 1.2 54.3 $\pm$ 0.8
SPD [140] 41.5 $\pm$ 0.5 44.3 $\pm$ 0.8 43.2 $\pm$ 1.2 46.2 $\pm$ 0.5 51.9 $\pm$ 0.5 55.1 $\pm$ 0.9
WS3D (Ours) 45.8 $\pm$ 0.5 49.2 $\pm$ 0.6 47.7 $\pm$ 0.9 53.9 $\pm$ 0.6 55.6 $\pm$ 0.7 58.2 $\pm$ 1.3
WS3D-Open (Ours) 47.2 $\pm$ 0.8 50.8 $\pm$ 0.6 51.1 $\pm$ 0.9 59.2 $\pm$ 0.8 59.0 $\pm$ 0.8 61.5 $\pm$ 1.6
WS3D++ (Ours) 55.3 $\pm$ 1.1 57.6 $\pm$ 1.2 57.9 $\pm$ 0.8 65.8 $\pm$ 1.2 64.6 $\pm$ 0.5 66.5 $\pm$ 1.1
TABLE IV: Comparisons of the open vocabulary learning performance. It can be demonstrated that our proposed approach provides very superior open-world recognition performance compared with the diverse SOTAs. WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
Datasets Models Few-shot settings
B15/N4 B12/N7 B10/N9
ScanNet [20] 3DGenZ [141] 20.6/56.0/12.6 19.8/35.5/13.3 12.0/63.6/6.6
3DTZSL [142] 10.5/36.7/6.1 3.8/36.6/2.0 7.8/55.5/4.2
LSeg3D [139] 0.0/64.4/0.0 0.9/55.7/0.1 1.8/68.4/0.9
PLA without caption [85] 39.7/68.3/28.0 24.5/70.0/14.8 25.7/75.6/15.5
PLA [85] 65.3/68.3/62.4 55.3/69.5/45.9 53.1/76.2/40.8
WS3D (Ours) 66.1/69.1/65.5 63.3/70.3/58.8 59.3/77.5/51.6
WS3D-Open (Ours) 68.3/70.6/67.1 64.2/71.4/59.9 59.7/77.9/51.6
WS3D++ (Ours) 72.3/69.1/73.3 73.3/70.0/63.6 65.2 / 78.9 / 58.8
Fully-supervised 74.5/68.4/79.1 73.6/72.0/72.8 69.9/75.8/64.9
Datasets Models Few-shot settings
B12/N3 B10/N5 B6/N9
NuScenes 3DGenZ [141] 01.6/53.3/00.8 01.9/44.6/01.0 01.1/52.6/00.5
3DTZSL [142] 01.2/21.0/0.6 06.4/17.1/03.9 2.61/18.52/03.15
LSeg-3D [85] 0.6/74.4/0.3 0.0/72.5/0.0 2.66/69.72/0.21
PLA without caption [85] 25.5/75.8/15.4 10.7/76.0/05.7 8.95/65.83/6.32
PLA [85] 47.7/73.4/35.4 24.3/73.1/14.5 15.63/60.32/12.38
WS3D (Ours) 51.3/76.6/30.9 41.5/72.5/20.8 35.2/40.9/29.8
WS3D-Open (Ours) 55.8/77.5/34.9 47.2/69.3/50.9 39.2/45.3/41.6
WS3D++ (Ours) 58.6/79.1/38.6 49.7/67.3/53.9 40.6/46.6/51.9
Fully-supervised 73.7/76.6/71.1 74.8/76.8/72.8 74.6/75.9/72.3
Figure 12: The WS3D++ corresponding detection and segmentation projections on six rendered 2D views through proposed multi-view rendering. It can be demonstrated that our proposed rendering has established an explicit modality association between the final 2D views and 3D views. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views. Best zoom-in for viewing.
@@
TABLE V: Comparison of the performance of instance segmentation, under various levels of supervision on ScanNet validation set. ’Sup-only-GPC’ denotes the model trained with only labeled data. ’WS3D’ denotes the model trained with our proposed methods. In the last row, we have shown the performance increase of WS3D. WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
@@
ScanNet Sup-only-GPC [121] 10.8 33.6 42.8 45.3 48.2 49.0 49.5 50.2 56.8
@@
WS3D (Ours) [4] 30.8 45.6 53.5 54.7 53.2 51.9 52.5 53.0 58.7
WS3D-Open (Ours) 31.2 50.6 55.8 57.9 56.8 58.7 58.9 59.5 65.1
WS3D++ (Ours) 39.2 54.8 57.3 58 +8.6 +9.7 +9.4 +9.3 +8.3

### v2

@@ 4.2 Finetuning Experimental Settings
Datasets. During the fine-tuning, to demonstrate the effectiveness for both data-efficient learning and open-world recognition of the proposed WS3D and WS3D++ under the limited scene reconstruction labeling scheme, we have tested it on various benchmarks, including S3DIS [19], ScanNet [102], and SemanticKITTI [16] for semantic segmentation, and ScanNet [102] for instance segmentation, respectively. Detailed information on each dataset and training details are put into the Appendix. The pre-training is conducted on ScanNet training set for indoor benchmark and on Waymo for outdoor benchmark.
@@
TABLE II: Boundary prediction average precision (AP) diverse label ratios for outdoor SemanticKITTI (SKITTI) benchmark and indoor ScanNet and S3DIS benchmark. Left of ’/’: The Average Precision (AP) of Semantic Boundary Prediction with focal Loss $L_{foc}$ Right of ’/’: The average precision with the cross entropy loss $L_{bce}$.
Labeling The APof Boundary Prediction (BP) WS3D++ Sem. Seg. mIOU%
Percentage SKITTI (Outdoor) ScanNet (Indoor) S3DIS (Indoor) SKITTI (Outdoor) ScanNet (Indoor) S3DIS (Indoor)
1% 56.7 / 61.6 55.8 / 59.1 56.2 / 61.7 51.6 58.8 55.8
5% 53.9 / 60.8 58.9 / 63.9 59.6 / 69.2 59.3 67.9 64.7
10% 55.7 / 61.5 62.2 / 65.7 62.6 / 66.9 68.7 76.7 68.2
15% 57.6 / 67.9 65.2 / 68.2 62.9 / 67.8 75.2 76.2 68.8
20% 59.7 / 66.5 66.3 / 69.8 65.2 / 68.7 72.9 79.8 75.8
25% 62.6 / 65.6 67.8 / 70.3 65.8 / 67.8 74.7 77.1 75.5
30% 64.6 / 70.2 71.6 / 73.8 68.9 / 72.8 78.7 76.3 75.9
40% 68.6 / 72.8 73.2 / 75.3 71.2 / 73.5 75.6 81.6 77.8
100% 83.7 / 88.7 76.9 / 82.6 76.8 / 83.7 83.9 88.8 82.9
TABLE III: Comparison of semantic segmentation results with different labeling percentages on ScanNet validation set, S3DIS validation set (Area 5), and SemanticKITTI validation set (Sequence 08). ’Sup-only-GPC’ denotes GPC model trained with only labeled data. ’WS3D’ denotes model trained with our proposed methods. We have shown the performance increase in the last row for each dataset, compared to merely trained models with labeled data (left value) and to the SOTAs GPC [103] (right value). The results in brackets (Pr) mean that the results after using our proposed pre-training approach to increase the final performance. It should be noted that the pre-training models for all weakly supervised approaches are the same with our proposed WS3D++ pre-training. It can be demonstrated that our proposed pre-training approach has a significant boost on the final scene parsing performance. Note that our subscript highlights the performance increment compared with the ”Sup-only-GPC”.
@@
ScanNet Sup-only-GPC–Baseline 40.9 48.1 57.2 61.3 64.0 65.3 67.1 68.8 72.9
Sup-only-GPC (Pr) 44.8 56.9 59.8 64.9 66.7 69.9 66.8 63.8 70.8
@@
Active-ST [107] (Pr) 49.9 59.5 67.9 67.9 75.3 70.6 73.7 75.2 78.1
WS3D++ 52.6 +11.7 59.1 +11.0 65.2 +8.0 66.9 +5.6 71.8 +6.6 72.0 +6.7 73.9 +5.5 75.3 +6.6 81.7 +8.8
WS3D++ (Pr) 58.8 +17.9 67.9 +19.8 76.5 +19.5 78.6 +14.9 79.8 +15.8 80.1 +11.8 80.8 +11.6 83.8 +12.8 88.8 +15.9
Merely Pre-training Stage 43.9 52.3 56.3 61.6 65.3 67.6 68.5 69.8 71.9
Sup-only-GPC–Baseline 35.3 44.5 52.9 53.8 59.9 60.3 61.2 62.6 66.4
S3DIS Sup-only-GPC (Pr) 48.6 51.7 60.8 61.9 70.8 72.8 69.7 68.8 73.6
@@
Mix3D [104] (Pr) 47.2 57.9 63.2 68.2 69.5 70.6 69.8 69.8 73.9
WS3D++ 48.6 +12.3 57.7 +12.7 61.2 +8.3 66.9 +11.6 70.6 +10.7 71.1 +10.8 72.6 +11.4 75.3 +12.7 82.0 +15.6
WS3D++ (Pr) 55.8 +19.5 64.7 +19.7 68.2 +15.3 68.8 +15.0 75.8 +15.9 75.5 +15.2 75.9 +14.7 77.8 +15.2 82.9 +16.5
Merely Pre-training Stage 45.2 52.8 58.7 63.8 68.6 69.8 71.2 72.6 73.9
SemanticKITTI Sup-only-GPC–Baseline 28.6 34.8 43.9 47.9 53.8 55.1 55.4 57.4 65.0
Sup-only-GPC (Pr) 36.3 41.8 50.3 55.6 57.6 63.8 60.8 59.8 67.8
@@
Mix3D [104] (Pr) 41.8 46.6 53.9 58.6 64.7 65.6 66.6 65.9 73.7
WS3D++ 46.8 +18.2 48.6 +13.8 55.2 +11.3 62.8 +14.9 65.9 +12.1 67.9 +12.8 68.6 +13.2 70.9 +13.5 76.8 +11.8
WS3D++ (Pr) 51.6 +33.0 59.3 +24.5 68.7 +24.8 69.8 +21.9 72.9 +19.1 74.7 +19.6 76.3 +20.9 76.6 +19.2 83.9 +18.9
Merely Pre-training Stage 42.9 46.6 51.7 61.7 63.8 67.8 69.6 71.8 75.8
@@
Implementation Details. For the task of semantic segmentation, we fine-tune the network for 500 epochs on a single NVIDIA 1080Ti GPU with a batch size of 16 during training. The initial learning rate is set to 1$\times$10-3 and is multiplied with 0.2 every 50 epochs. We implement it in PyTorch and optimize it with Adam optimizer [109]. We set the hyperparameter $\gamma$ as 0.8 to ensure that merely highly confident prediction can be used for network optimization. $\epsilon$ is set to 0.5. We empirically choose $\alpha=\beta=1$, while $\lambda_{u}=0.1$. For instance segmentation, we train the network for 580 epochs on a single NVIDIA 1080Ti GPU with a batch size of 8 during training. The other settings are the same as the semantic segmentation task.
@@ 4.3 Data-efficient 3D Semantic Segmentation
Overall Experimental Results. For the semantic segmentation, we have tested WS3D++ on versatile indoor and outdoor benchmarks, including ScanNet [18], S3DIS [19], and SemanticKITTI [16]. We have done extensive experiments with limited labeled data, e.g., only {1%, 5%, 10%, 15%, 20%, 25%, 30%, 40%, 100%} data in training set are available as labeled data. The qualitative results are shown in Fig. 5. In the meanwhile, the quantitative semantic segmentation performance is summarized in Table III. As mentioned, we have used the voxel-based method SparseConv [59] as the backbone. Our WSL model significantly surpasses the supervised-only model in GPC that is merely trained with labeled data, showing that our WSL can effectively make use of the unlabeled data to enhance the feature discrimination capacity of the model. Also, it can be observed the increment of performance is more obvious when the unlabeled data percentage is larger. For example, the performance increase on SemanticKITTI is 10.3% for the 1% labeling percentage, 5.8% for the 40% labeling percentage, and 1.9% for the 100% labeling percentage. This can be possibly explained by the fact that for more unlabeled data, our proposed WS3D++ can extract more meaningful semantic information from the unlabeled data based on our boundary-guided energy-based loss and confidence-guided region-level contrastive learning design. In addition, compared with current SOTA GPC, our proposed WS3D++ also achieves consistently better results in semantic segmentation performance, especially when faced with very limited label circumstances (e.g. 1% labeling points). In that case, WS3D++ outperforms GPC by 3.3%, 7.1%, and 4.2% for ScanNet, S3DIS, and SemanticKITTI, respectively. Fig. 5 shows that we can provide comparable performance compared with fully supervised SOTAs BAAF-Net [35] and Cylinder3D [34] on SemanticKITTI with 5% labels. As shown in Table III, the performance of our enhanced approach WS3D++ has remarkably increased performance compared with WS3D++ and previous SOTAs, demonstrating the effectiveness of our proposed vision-language knowledge-associated pre-training.
We have reported the experimental results about semantic boundary prediction as demonstrated in Table 1 of our revised manuscript, it is demonstrated that our proposed approach can provide a very high-quality label at a very low-label regime, and the boundary prediction average precision can be kept at more than 50% (50.9% at least). The phenomenon has also been found at previous research including SQN, Lasermix, and our Weaklab3DNet [110]. The reason behind is that: As point clouds are essentially samples of the 3D world, the distribution of points in a very close local neighborhood is comparatively homogeneous, revealing strong semantic similarity/homogeneity. Moreover, our proposed weakly supervised approach can be regarded as an amplification of those rather sparse supervision signals, which largely facilitates ultimate semantic boundary prediction. As we have demonstrated in Table I, the average precision (AP) of boundary prediction can still be maintained at a relatively high value (more than 50.9%). Our proposed approach can realize promising performance on outdoor benchmark SemanticKITTI and indoor benchmarks including S3DIS and ScanNet, which demonstrates the superior effectiveness of our proposed weakly supervised semantic boundary prediction.
Figure 13: The WS3D++ corresponding detection and segmentation projections on six rendered 2D views through our proposed multi-view rendering approach. It can be demonstrated that our proposed rendering has established an explicit modality association between the final 2D views and 3D views. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views. Best zoom-in for viewing.
@@
TABLE IV: Comparison of experimental results on 20% and fully labeled case for the task of inductive and transductive learning for our proposed WS3D++, respectively. In transductive learning, the test set is also utilized for network training. We test the task of semantic segmentation on ScanNet, S3DIS, and SemanticKITTI with the evaluation metric of mIoU(%).
@@
ScanNetv2-WS3D++ 69.9 71.8 85.8 78.9 81.7 87.8
ScanNetv2-WS3D++ (Pr) 75.3 79.8 87.8 85.6 88.8 91.8
S3DIS WS3D++ 62.8 70.6 75.8 69.7 82.0 86.8
S3DIS WS3D++ (Pr) 68.7 75.8 78.7 73.8 82.9 87.3
Semantic KITTI Val. WS3D++ 58.9 65.9 79.3 74.3 76.8 88.8
Semantic KITTI Val. WS3D++ (Pr) 62.8 72.9 69.1 68.6 83.9 85.5
Semantic KITTI Test. WS3D++ 66.7 71.8 72.6 71.6 77.9 79.6
Semantic KITTI Test. WS3D++ (Pr) 70.8 75.9 78.8 79.1 84.3 86.5
@@
TABLE V: Comparison of current State-of-the-art (SOTA) approaches in the limited reconstruction case for the 3D object detection tasks with different ratios of labeled data. The mean average precision (mAP) given also with the mean $\pm$ standard deviation in three runs of diverse random splits are reported. Baseline is the circumstance which is trained with merely labelled data. It can be demonstrated our proposed pre-training approach have improved the performance of 3D object detection by a considerable margin.
Datasets Models 5% 10% 20% 100%
Induct. Transduct. Induct. Transduct. Induct. Transduct. Induct. Transduct.
SUN RGB-D [111, 110] Baseline 29.9 $\pm$ 1.5 33.5 $\pm$ 0.8 34.4 $\pm$ 1.1 40.7 $\pm$ 0.9 41.1 $\pm$ 0.3 47.5 $\pm$ 0.5 51.7 $\pm$ 0.9 53.8 $\pm$ 0.7
Merely the Pre-training Stage 35.8 $\pm$ 1.1 38.2 $\pm$ 1.1 39.3 $\pm$ 0.7 45.1 $\pm$ 0.7 46.1 $\pm$ 0.9 56.9 $\pm$ 0.7 51.9 $\pm$ 1.1 65.8 $\pm$ 1.5
SESS [112] 34.2 $\pm$ 2.0 38.1 $\pm$ 0.7 42.9 $\pm$ 0.8 45.3$\pm$ 0.9 47.9 $\pm$ 0.4 51.6 $\pm$ 0.2 50.7 $\pm$ 0.5 53.7 $\pm$ 0.5
SESS (Pr) [112] 42.8 $\pm$ 3.2 42.6 $\pm$ 0.7 48.2 $\pm$ 0.8 53.8 $\pm$ 0.9 50.5 $\pm$ 0.6 56.3 $\pm$ 0.7 55.2 $\pm$ 0.3 63.9 $\pm$ 0.6
3D-IOUMatch [92] 39.1 $\pm$ 1.9 46.3 $\pm$ 0.7 45.5 $\pm$ 1.5 53.5 $\pm$ 0.3 51.6 $\pm$ 0.5 55.6 $\pm$ 0.8 55.6 $\pm$ 0.3 62.3 $\pm$ 0.7
3D-IOUMatch (Pr) [92] 43.6 $\pm$ 3.2 48.5 $\pm$ 0.9 48.5 $\pm$ 1.8 56.9 $\pm$ 0.5 59.8 $\pm$ 0.7 56.8 $\pm$ 0.8 65.7 $\pm$ 0.5 69.1 $\pm$ 0.7
SPD [113] 38.5 $\pm$ 0.7 44.3 $\pm$ 0.8 46.0 $\pm$ 1.0 51.7 $\pm$ 0.9 60.9 $\pm$ 0.5 56.9 $\pm$ 0.9 67.9 $\pm$ 0.7 73.9 $\pm$ 0.8
SPD [113] (Pr) 43.6 $\pm$ 1.2 47.8 $\pm$ 1.7 49.6 $\pm$ 0.8 53.8 $\pm$ 1.1 58.6 $\pm$ 1.6 68.9 $\pm$ 2.5 69.2 $\pm$ 0.9 83.8 $\pm$ 0.5
WS3D++ (Ours) 52.7 $\pm$ 0.7 54.5 $\pm$ 0.8 53.7 $\pm$ 0.9 63.8 $\pm$ 0.8 69.8 $\pm$ 0.7 72.9 $\pm$ 1.1 68.6 $\pm$ 0.9 82.8 $\pm$ 0.5
WS3D++ (Ours) (Pr) 56.9 $\pm$ 0.8 59.7 $\pm$ 1.5 58.7 $\pm$ 1.8 68.9 $\pm$ 1.3 68.6 $\pm$ 1.8 73.7 $\pm$ 1.6 72.6 $\pm$ 1.1 83.9 $\pm$ 0.9
ScanNet-V2 [18] Baseline 27.9 $\pm$ 0.5 30.8 $\pm$ 1.5 31.1 $\pm$ 0.7 33.2 $\pm$ 0.5 41.6 $\pm$ 0.9 44.5 $\pm$ 1.2 45.8 $\pm$ 0.8 48.9 $\pm$ 0.9
Merely the Pre-training Stage 35.2 $\pm$ 1.5 32.8 $\pm$ 1.1 35.6 $\pm$ 0.8 36.6 $\pm$ 0.6 43.9 $\pm$ 1.1 47.9 $\pm$ 0.8 49.3 $\pm$ 0.7 56.8 $\pm$ 0.9
SESS [112] 32.2 $\pm$ 0.8 36.8 $\pm$ 1.1 39.7 $\pm$ 0.9 44.5 $\pm$ 0.5 48.2 $\pm$ 0.4 49.6 $\pm$ 0.7 55.8 $\pm$ 0.1 56.8 $\pm$ 0.7
SESS [112] (Pr) 35.9 $\pm$ 0.9 43.9 $\pm$ 1.2 43.5 $\pm$ 1.2 48.9 $\pm$ 0.5 49.6 $\pm$ 0.6 53.2 $\pm$ 0.7 56.9 $\pm$ 0.6 59.6 $\pm$ 0.9
3D-IOUMatch [92] 40.0 $\pm$ 0.9 46.3 $\pm$ 0.7 47.2 $\pm$ 0.4 49.6 $\pm$ 0.6 52.8 $\pm$ 1.2 62.9 $\pm$ 0.8 56.9 $\pm$ 1.2 63.7 $\pm$ 0.8
3D-IOUMatch (Pr) [92] 45.6 $\pm$ 2.0 52.8 $\pm$ 0.8 53.3 $\pm$ 1.5 55.6 $\pm$ 1.7 57.7 $\pm$ 1.2 59.6 $\pm$ 0.8 61.8 $\pm$ 1.6 66.9 $\pm$ 1.3
SPD [113] 41.5 $\pm$ 0.5 44.3 $\pm$ 0.8 43.2 $\pm$ 1.2 46.2 $\pm$ 0.5 51.9 $\pm$ 0.5 55.6 $\pm$ 0.9 54.8 $\pm$ 0.5 58.6 $\pm$ 0.9
SPD [113] (Pr) 45.2 $\pm$ 0.7 48.8 $\pm$ 1.9 52.8 $\pm$ 1.6 55.2 $\pm$ 0.7 53.8 $\pm$ 1.8 65.2 $\pm$ 2.5 59.9 $\pm$ 0.5 68.9 $\pm$ 1.6
WS3D++ (Ours) 55.3 $\pm$ 1.1 57.6 $\pm$ 1.2 59.8 $\pm$ 0.8 66.8 $\pm$ 1.2 64.6 $\pm$ 0.5 72.8 $\pm$ 1.1 68.7 $\pm$ 0.5 76.6 $\pm$ 1.1
WS3D++ (Ours) (Pr) 66.6 $\pm$ 1.5 68.1 $\pm$ 1.7 67.6 $\pm$ 1.6 74.5 $\pm$ 1.3 69.5 $\pm$ 0.9 76.9 $\pm$ 0.9 73.6 $\pm$ 0.7 77.8 $\pm$ 0.6
@@
TABLE VI: Comparison of the performance of instance segmentation, under various levels of supervision on ScanNet validation set. ’Sup-only-GPC’ denotes the model trained with only labeled data. ’WS3D’ denotes the model trained with our proposed methods. In the last row, we have shown the performance increase of WS3D. WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [7] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$. WS3D-Open is examined to validate the effect of hierarchical feature-aligned pretraining in improving data efficiency. The results of our proposed approach are highlighted in orange.
@@
ScanNet Merely the Pre-training Stage 39.7 46.7 53.8 55.6 58.6 63.5 65.2 65.7 65.8
Sup-only-GPC–Baseline [103] 10.8 33.6 42.8 45.3 48.2 49.0 49.5 50.2 56.8
Sup-only-GPC (Pr) [103] 16.8 35.9 33.8 48.9 56.6 57.9 58.8 66.8 68.9
@@
SPIB_Ins (Pr) [114] 23.7 52.8 52.9 55.3 54.8 55.8 56.9 58.8 64.8
WS3D++ (Ours) 45.9 61.8 66.8 67.7 65.1 65.2 62.7 68.7 58.8
WS3D++ (Ours) (Pr) 48.2 72.2 71.6 70.5 67.9 68.7 71.6 75.8 73.9